Find the amplitude, period, and phase shift.
y = 10*cos(½x)

Answer by Edwin McCravy(20060)

Find the amplitude, period, and phase shift.
y = 10*cos(½x)

Rule:

For either y = A*sin(Bx+C) or y = A*cos(Bx+C)

Amplitude = |A|
Period = 2p/B
Phase shift = C/B and if that is positive the phase 
shift is left and if negative the phase shift is right.

y = 10*cos(½x)

y = 10*cos(½x+0)

A = 10, B = ½, C = 0.

Amplitude = |A| = |10| = 10
Period = 2p/B = 2p/½ = 2p*2 = 4p
Phase shift = C/B = 0/½ = 0 (No phase shift).



This is the graph of one period of y = 10*cos(½x).
Notice that it goes from 0 to 4p.
Notice that the x-intercepts are at 0, p,
2p, 3p, and 4p 

Notice that the three green lines are 10 units each,
which is the amplitude.

Edwin